# Comparing Mass Mapping Reconstruction Methods with Minkowski Functionals

### v1

@@ 1 Introduction
Various non-parametric mass mapping reconstruction methods have been developed to address this problem and take into account noise and masks in the data to solve the inverse problem. Linear approaches like Kaiser Squires (KS) (Kaiser & Squires, 1993), Wiener filter (WF) (Lahav et al., 1994), Simon et al. (2009), and VanderPlas et al. (2011) and non-linear approaches like GLIMPSE (Leonard et al., 2014), Glimpse2D (Lanusse et al., 2016), KS+ (Pires et al., 2020), DeepMass (Jeffrey et al., 2020), MCAlens (Starck et al., 2021), DarkMappy (Price et al., 2021), KaRMMa (Fiedorowicz et al., 2022), DeepPosterior (Remy et al., 2022), SKS+ (Kansal, 2023), and others (Marshall et al., 2002; Simon et al., 2011b; Leonard et al., 2012; Alsing et al., 2015; Schneider et al., 2017; Horowitz et al., 2019; Porqueres et al., 2021) have all been shown to be effective. The original method, Kaiser-Squires (KS) is a linear inversion of the shear map into a convergence map, but it cannot take noise or boundary effects into account (Kaiser & Squires, 1993). Wiener filtering is also a linear filter, but unlike KS, it can transform noisy data well (Lahav et al., 1994). Sparsity methods like GLIPMSE and DarkMappy are non-linear and able to reconstruct information in the non-Gaussian regime (Leonard et al., 2014; Price et al., 2021). Some methods like DarkMappy and DeepPosterior are able to approximate the uncertainty that comes from reconstruction. Other methods including DeepPosterior have employed machine learning as a useful tool in denoising convergence signal (Jeffrey et al., 2020; Shirasaki et al., 2021; Remy et al., 2022). In this paper we explore four reconstruction methods: Kaiser Squires, Wiener filter, DarkMappy, and DeepMass.
@@
In this paper we use MFs as a test of different convergence reconstruction methods, using them to quantify the information (particularly on small scales) that the methods can reconstruct in maps. We compare the four different mass mapping methods KS, WF, DarkMappy, and DeepMass, applying them to simulated images and using MFs to constrain CDM parameters and thus computing figures of merit to describe method power.
@@ 2.3 Wiener filter
A Wiener filter (WF) is a linear inversion of signal and noise together, aiming to suppress the latter. The WF method uses the expected power spectrum of the convergence field at a fiducial cosmology and of the noise (Simon et al., 2011a). In Fourier space:
where S is the covariance matrix of power spectrum signal and N is the noise covariance matrix (Lahav et al., 1994; Zaroubi et al., 1995). However, in filtering out the noise, WF smooths out small-scales features in the map more generally. While WF can reconstruct the field with precision on large scales, it does not typically retain small-scale information in convergence maps as well (Simon et al., 2011a; Horowitz et al., 2019).
@@ 2.6 Minkowski functionals
Minkowski functionals are field integrals that characterise the topological properties of continuous fields (Minkowski, 1903; Zürcher et al., 2021). In this work we measure these mathematical descriptors from a convergence field. MFs measure the properties of excursion sets of this field, which are given by the region of the field above a given threshold (Parroni et al., 2020). The first three functionals quantify area, perimeter, and mean curvature of an excursion set and are defined as
@@ 3.1.2 Noise and Mask
At our chosen resolution our maps are in the low noise regime. We only take shape noise into account, disregarding any other observational effects. To obtain the noise standard deviation per pixel, we use
where the shape dispersion is 0.26 per galaxy, is a DES Y3-like number density (10 galaxies/arcmin), is the four DES Y3-like redshift bins, and is the pixel size in arcmin. With 256 total pixels and 15 15 square degree maps, the pixel size is 3.5 arcmin and the noise level is 0.047 per pixel. The map noise is computed as an uncorrelated random Gaussian field with mean 0 and from Eq. (22).
We apply a mask to our maps by hand to approximate the structure of real masks, covering approximately 2.8% of the map, as seen in Figure 2.
The noise and mask are added at the shear step. A Gaussian kernel with a standard deviation of one pixel has to be applied to the masked shear map before reconstruction. This smoothing fills in gaps and decreases the resolution of the maps. This is a crucial step for Kaiser Squires, which only works for continuous fields.
@@ 3.1.3 Applications of Reconstruction Methods
The noise matrix in Eq. (9) is a uniform matrix with value from Eq. (22). It describes the statistics of the noise in the convergence map.
@@ 5 Conclusion
This difference aligns closely with observed variations in average MF DeepMass outputs from data trained on the correct or incorrect model. Both effects are comparable with the change in cosmology for . These findings demonstrate a strong dependence of DeepMass on the cosmology of the training dataset, emphasising the need for the consideration of input cosmology in model training to ensure accurate and reliable results.
The remaining methods, KS, WF, and DarkMappy are comparable at the scales we use here. DarkMappy was expected to outperform KS and WF. One reason it did not could be that it was designed for cluster analysis and may be tuned for such scenarios rather than our wide-field tests; alternative choices of its parameters and wavelet dictionary could improve its performance. Another could be the resolution of the field. In this paper we have used CosmoGridV1 simulations with DES Y3-like redshift bins and noise, which have led to a more limited resolution. We expect an improvement in resolution with less noisy survey conditions like for LSST or in future LSST-like forecasts, as well as with newer simulations.
@@
We use Minkowski functionals to evaluate the different reconstruction methods, but there are many other statistics that could be used. Other higher order statistics may yield different results and show more distinction between the contours. Likewise, we have chosen four reconstruction methods here, representative of the different approaches to mass mapping. Other methods may be as or more precise.

### v2

@@ 1 Introduction
Various non-parametric mass mapping reconstruction methods have been developed to address this problem and take into account noise and masks in the data to solve the inverse problem. Linear approaches like Kaiser Squires (KS) (Kaiser & Squires, 1993), Wiener filter (WF) (Lahav et al., 1994), Simon et al. (2009), and VanderPlas et al. (2011) and non-linear approaches like GLIMPSE (Leonard et al., 2014), Glimpse2D (Lanusse et al., 2016), KS+ (Pires et al., 2020), DeepMass (Jeffrey et al., 2020), MCAlens (Starck et al., 2021), DarkMappy (Price et al., 2021), KaRMMa (Fiedorowicz et al., 2022), DeepPosterior (Remy et al., 2022), SKS+ (Kansal, 2023), and others (Marshall et al., 2002; Simon et al., 2011; Leonard et al., 2012; Alsing et al., 2015; Schneider et al., 2017; Horowitz et al., 2019; Porqueres et al., 2021) have all been shown to be effective. The original method, Kaiser-Squires (KS) is a linear inversion of the shear map into a convergence map, but it cannot take noise or boundary effects into account (Kaiser & Squires, 1993). Wiener filtering is also a linear filter, but unlike KS, it can transform noisy data well (Lahav et al., 1994). Sparsity methods like GLIPMSE and DarkMappy are non-linear and able to reconstruct information in the non-Gaussian regime (Leonard et al., 2014; Price et al., 2021). Some methods like DarkMappy and DeepPosterior are able to approximate the uncertainty that comes from reconstruction. Other methods including DeepPosterior have employed machine learning as a useful tool in denoising convergence signal (Jeffrey et al., 2020; Shirasaki et al., 2021; Remy et al., 2022). In this paper we explore four reconstruction methods: Kaiser Squires, Wiener filter, DarkMappy, and DeepMass.
@@
In this paper, we use MFs as a test of different convergence reconstruction methods, using them to quantify the information (particularly on small scales) that the methods can reconstruct in maps. We compare the four different mass mapping methods KS, WF, DarkMappy, and DeepMass, applying them to simulated images and using MFs to constrain CDM parameters and thus computing figures of merit to describe method power.
@@ 2.3 Wiener filter
A Wiener filter (WF) is a linear inversion of signal and noise together, aiming to suppress the latter. The WF method uses the expected power spectrum of the convergence field at a fiducial cosmology and of the noise (Simon et al., 2011). In Fourier space:
where S is the covariance matrix of power spectrum signal and N is the noise covariance matrix (Lahav et al., 1994; Zaroubi et al., 1995). However, in filtering out the noise, WF smooths out small-scales features in the map more generally. While WF can reconstruct the field with precision on large scales, it does not typically retain small-scale information in convergence maps as well (Simon et al., 2011; Horowitz et al., 2019).
WF balances noise reduction and signal preservation. Underestimating the noise level causes WF to misinterpret noise as signal, resulting in a loss of accuracy; overestimating the noise causes WF to suppress signal as noise, affecting the reconstruction precision as the noise is not reduced properly (Simon et al., 2011; Horowitz et al., 2019).
@@ 2.6 Minkowski functionals
Minkowski functionals are field integrals that characterise the topological properties of continuous fields (Minkowski, 1903; Zürcher et al., 2021). In this work we measure these mathematical descriptors from a convergence field. MFs measure the properties of excursion sets of this field, which are given by the region of the field above a given threshold (Parroni et al., 2020). The first three functionals quantify area, perimeter, and mean curvature of an excursion set and can be written777We write curvature integral summary statistics of MFs in Eqs. (17), (18), and (19), which are valid for a field with no boundary. See Mecke et al. (1993) for the most general definitions. as
@@ 3.1.2 Noise and Mask
At our chosen resolution, our maps are in the low noise regime. We only take shape noise into account, disregarding any other observational effects. To obtain the noise standard deviation per pixel, we use
where the shape dispersion is 0.26 per galaxy, is a DES Y3-like number density (10 galaxies/arcmin2), is the four DES Y3-like redshift bins, and is the pixel size in arcmin. With 2562 total pixels and 15 15 square degree maps, the pixel size is 3.5 arcmin and the noise level is 0.047 per pixel. The map noise is computed as an uncorrelated random Gaussian field with mean 0 and from Eq. (22).
We apply a mask to our maps by hand to approximate the structure of real masks, covering approximately 2.8% of the map, as seen in Figure 2. The regions in the mask do not correspond to specific simulations of real physical processes or artefacts, but rather we have manually created a mock mask matching the general shape of typical DES masks (Gatti et al., 2021a).
After applying an inverse Kaiser-Squires transform to the CosmoGridV1 convergence maps, we add the noise and mask to the resulting shear maps. A Gaussian kernel with a standard deviation of one pixel has to be applied to the masked shear map before reconstruction. To omit the masked pixels, we compute
where is the input map with the masked pixels set to zero and is a binary mask containing ones for unmasked pixels. This smoothing fills in gaps and decreases the resolution of the maps. This is a crucial step for Kaiser Squires, which only works for continuous fields.
@@ 3.1.3 Applications of Reconstruction Methods
The noise covariance matrix in Eq. (9) is a diagonal matrix with value from Eq. (22). It describes the statistics of the noise in the convergence map.
@@ 3.4 Figures of Merit
Our code used to transform a Fisher matrix into a contour for a given set of cosmological parameters (including and ) is available here.888https://github.com/nishagrewal/fisher_to_contour
@@ 5 Conclusion
As DeepMass denoises the Wiener filter (WF) output convergence, the smoothing in WF implies a certain scale where the signal dominates for DeepMass. DeepMass has the capability to go beyond this with either higher resolution maps, lower noise, or a different initialisation that does not rely on WF.
@@
This difference aligns closely with observed variations in average MF DeepMass outputs from data trained on the correct or incorrect model. Both effects are comparable with the change in cosmology for . These findings demonstrate a strong dependence of DeepMass on the cosmology of the training dataset, emphasising the need for the consideration of input cosmology in model training to ensure accurate and reliable results. While a larger training set could improve the signal reconstruction, a simulation-based inference framework is perhaps a safer solution since it can more explicitly incorporate these uncertainties.
The remaining methods, KS, WF, and DarkMappy are comparable at the scales we use here. DarkMappy was expected to outperform KS and WF. One reason it did not could be that it was designed for cluster analysis and may be tuned for such scenarios rather than our wide-field tests; alternative choices of its parameters and wavelet dictionary could improve its performance. Another could be the resolution of the field. In this paper, we have used CosmoGridV1 simulations with DES Y3-like redshift bins and noise, which have led to a more limited resolution. We expect an improvement in resolution with less noisy survey conditions like for LSST or in future LSST-like forecasts, as well as with newer simulations.
@@
We use Minkowski functionals to evaluate the different reconstruction methods, as they are quick to measure, unbiased, and typically resilient to systematic uncertainties. However, there are many other (combinations of) statistics that could be used to probe the underlying dark matter distribution. The two-point correlation function measures two-point information, and other higher order statistics may yield different results and show more distinction between the contours. Likewise, we have chosen four reconstruction methods here, representative of the different approaches to mass mapping. Other methods may be as or more precise.